OpenStudy (anonymous):

-(10)^-1 A.-1/10 B.-1/-1^10 C.1/10 D.10

jhonyy9 (jhonyy9):

use x^(-1) = 1/x

jhonyy9 (jhonyy9):

but -x^(-1) = - 1/x ok. ?

OpenStudy (anonymous):

ok thx

jhonyy9 (jhonyy9):

np was my pleasure